# Integrating Mechanistic Modeling and Machine Learning to Study CD4+/CD8+ CAR-T Cell Dynamics with Tumor Antigen Regulation

### v1

@@ Abstract
Chimeric antigen receptor (CAR) T cell therapy has shown remarkable success in hematological malignancies, yet patient responses remain highly variable and the roles of CD4+ and CD8+ subsets are not fully understood. We present an extended mathematical framework of CAR-T cell dynamics that explicitly models CD4+ helper and CD8+ cytotoxic lineages and their interactions with tumor antigen burden. Building on the Kirouac et al. (2023) model of antigen-regulated memory–effector–exhaustion transitions, our system of differential equations incorporates cytokine-mediated modulation of CD8+ proliferation, cytotoxicity, and memory regeneration by CD4+ T cells. Sensitivity analyses identify effector proliferation burst size, antigen turnover, and CD8+ expansion rates as dominant determinants of treatment outcome. Virtual patient simulations reproduce clinical findings that a 1:1 CD4:CD8 ratio enhances CAR-T expansion and tumor clearance relative to CD8-only products. Finally, we integrate a feed-forward neural network trained on noisy virtual patient data to improve predictive robustness, and apply SHAP analysis to interpret the network’s predictions and compare them with mechanistic sensitivity analyses. This work highlights the synergistic roles of CD4+ and CD8+ CAR-T cells and provides a quantitative foundation for optimizing treatments and patient stratification.
@@ 1 Introduction
CD4+ and CD8+ T cells are distinct lineages of T cells (Germain et al. 2002), each with unique and complementary roles in the immune response. CD8+ T cells, known as "cytotoxic" T cells, specialize in directly killing foreign cells (Anderson et al. 2006). They are equipped with the ability to recognize antigens presented by infected or malignant cells and trigger apoptosis (cell death) in those cells. CD4+ T cells, often referred to as "helper" T cells, play a crucial role in orchestrating the immune response by secreting cytokines that activate other immune cells, including CD8+ T cells (Ahrends et al. 2013; Ahrends et al. 2017). Importantly, CD4+ and CD8+ T cells are separate lineages; they do not switch from one to the other, but instead, each lineage maintains its specialized function throughout its lifespan. In cancer immunotherapy, the roles of both CD4+ and CD8+ T cells are critical, as CD4+ cells provide necessary signals to optimize CD8+ T cell activity, creating a synergistic effect in tumor elimination (Boulch et al. 2021).
It has been shown that a 1:1 ratio of CD4+ to CD8+ CAR-T cells yields optimal therapeutic outcomes (as opposed to an unspecified ratio dependent on patient T cell composition), balancing efficacy and toxicity across multiple dosing levels (Turtle et al. 2016). Motivated by these findings, we extend the Kirouac et al. model to explicitly include CD4+ and CD8+ T cell lineages and their distinct functional roles. This formulation enables the simulation of various CD4:CD8 ratios and their impact on tumor–immune dynamics. We extend this framework by modeling separate CD4+ and CD8+ CAR-T cell lineages, thereby enabling the investigation of their distinct functional roles and synergistic interactions (Figure 1).
@@ 2 Methods
In this study, we introduce a coupled mathematical model to simulate the dynamics of CD4+ and CD8+ CAR-T cells in response to tumor antigen regulation. The model captures the antigen-driven transitions of T cells between memory, effector, and exhausted states as proposed in Kirouac et al. (2023), and incorporates the modulation of CD8+ CAR-T cells by CD4+ CAR-T cells. By modeling the interactions between these cell lineages and tumor cells, we aim to predict the therapeutic outcomes of CAR-T cell therapies with varying CD4:CD8 ratios. The model also allows for an exploration of the effects of patient-specific variations on treatment response. A schematic representation of the model is shown in Figure 1.
@@ 2.1 Tumor and Antigen Dynamics
$\frac{dB}{dt}$ $=\mu_{B}\left(1-\frac{B}{B_{\max}}\right)B-k_{\text{kill}}^{CD8}\left(1+\eta_{k}\,\frac{T_{E2}^{CD4}}{K_{\eta}+T_{E2}^{CD4}}\right)\,\left(\frac{{T_{E2}^{CD8}}^{k_{t}}}{TK50^{k_{t}}+{T_{E2}^{CD8}}^{k_{t}}}\right)B$
@@
where $\eta_{k}$ denotes the maximum fold-increase in CD8+ cytotoxicity attributable to CD4+ T cell-derived cytokines, and $K_{\eta}$ is the half-saturation constant (Ahrends et al. 2017). Biologically, this formulation reflects the role of CD4+ T cells, particularly Th1-like CAR-T cells, in secreting cytokines such as interferon-gamma (IFN-$\gamma$) and interleukin-2 (IL-2), which are known to enhance the activation, survival, and cytotoxic function of CD8+ T cells (Boulch et al. 2021). IFN-$\gamma$ also promotes tumor antigen presentation and induces tumor cell apoptosis, thereby indirectly causing cytotoxicity (Bove et al. 2023); this is reflected in the final term of Eq. (1).
@@ 2.2 CD8+ CAR-T Cell Dynamics
The dynamics of CD8+ CAR-T cells are modelled by a set of differential equations that model the transitions between different T cell states: memory ($T_{M}^{CD8}$), early effector ($T_{E1}^{CD8}$), late effector ($T_{E2}^{CD8}$), and exhausted ($T_{X}^{CD8}$) (Kirouac et al. 2023). These equations take into account the proliferation, differentiation, death rates, and interactions with tumor antigens ($B_{A}$).
@@
$\frac{dT_{M}^{CD8}}{dt}$ $=\mu_{M}^{CD8}\left(2f_{\text{max}}^{CD8}\left(1-\frac{B_{A}^{k_{m}^{CD8}}}{B50^{CD8^{k_{m}}}+B_{A}^{k_{m}^{CD8}}}\right)-1\right)T_{M}^{CD8}$
@@ 2.3 CD4+ CAR-T Cell Dynamics
CD4+ T cells, or helper T cells, play an essential role in coordinating the immune response. Unlike CD8+ T cells, the main role of CD4+ cells is not directly killing infected or cancerous cells (although they can induce cell death by themselves) (Boulch et al. 2021). Instead, their main function involves enhancing the function of other immune cells, including CD8+ T cells, through the secretion of cytokines such as interferon-gamma (IFN-$\gamma$) and interleukin-2 (IL-2) (Ahrends et al. 2017). These cytokines stimulate the expansion and survival of CD8+ T cells, thereby augmenting their cytotoxic effects on tumor cells. In CAR-T therapy, CD4+ T cells are critical for sustaining an effective immune response, promoting memory formation, and improving tumor clearance (Boulch et al. 2021).
@@
$\frac{dT_{M}^{CD4}}{dt}$ $=\mu_{M}^{CD4}\left(2f_{\text{max}}^{CD4}\left(1-\frac{B_{A}^{k_{m}^{CD4}}}{B50^{CD4^{k_{m}}}+B_{A}^{k_{m}^{CD4}}}\right)-1\right)T_{M}^{CD4}$
@@ 2.4 Model Parameters
A list of parameters used in the model is given in the Supplementary Materials (Table 1). Kirouac et al. reported physiologically feasible ranges and averages for the parameters in their original model, which did not have separate CD4+ and CD8+ compartments (Kirouac et al. 2023). However, several parameters are likely to differ biologically between these two lineages (Table 2). CD8+ CAR-T cells are primarily responsible for direct cytotoxicity and tend to differentiate more rapidly into effector cells ($\mu_{E}$) but are also more prone to exhaustion ($k_{ex}$) and activation-induced cell death ($d_{E}$, $d_{X}$) (Turtle et al. 2016; Hinrichs et al. 2011; Foulds et al. 2002; Fraietta et al. 2018; Gauthier et al. 2021; Blank et al. 2019; Siddiqui et al. 2019). In contrast, CD4+ CAR-T cells play a predominantly helper role through cytokine production (e.g., IFN-$\gamma$, IL-2), are generally more resistant to exhaustion, and may exhibit greater longevity and plasticity, potentially resulting in higher rates of reversion to memory ($r_{M}$). As such, when fitting the model to data or simulating lineage-specific behavior, these parameters should be relaxed and estimated separately for CD4+ and CD8+ populations. This refinement is supported by experimental evidence from preclinical and clinical studies, including those by Turtle et al. (2016), Fraietta et al. (2018), and Gauthier et al. (2021), which consistently demonstrate phenotypic and functional distinctions between CD4+ and CD8+ CAR-T cells in vivo. .
@@ 2.5 Sensitivity Analysis
There are two main types of sensitivity analysis: local and global. Local sensitivity analysis involves perturbing each parameter individually from a baseline value while holding other parameters constant, and looking at the change in output (Qian et al. 2020). In this context, we applied finite difference perturbations (1%) to each parameter (starting from average patient parameter values) and initial condition individually while tracking changes in the area under the curve (AUC) of tumor burden over 1 year. Sensitivities were computed as normalized relative changes, defined as the relative change in the output (area under the curve) normalized by the relative change in the input, allowing us to rank the most influential parameters under the baseline scenario.
@@ 3.1 Sensitivity Analysis
Local sensitivity analysis (Figure 2) revealed that the most sensitive parameters were the proliferation burst factor during antigen stimulation ($N$), tumor antigen generation and clearance rates ($k_{b1}$ and $k_{b2}$), CD8+ T cell expansion dynamics (including the effector proliferation rate $k_{e}^{CD8}$ and proliferation rate $u_{m}^{CD8}$), and tumor-related parameters such as the antigen half-max ($B_{50}^{CD8}$) (Figure 2). These parameters exhibited the highest relative sensitivities; that is, small changes in their values can lead to substantial differences in tumor burden dynamics within the model.
@@
Global sensitivity analysis (Figures 3, 4) found that the parameter representing the number of effector cell doublings ($N$) exhibits the highest sensitivity in our model, with both PRCC and Sobol analyses indicating it as the strongest driver of tumor burden dynamics. Parameters governing the dynamics of the antigen-bound complex ($k_{b1}$, $k_{b2}$) and the CD8+ effector proliferation rate ($k_{e}^{CD8}$) also showed high sensitivities. These findings align with prior analysis by Kirouac et al. of the original mechanistic framework (2023), which identified effector proliferation and memory regeneration as critical determinants of treatment response. Our analyses suggest that modulating parameters such as $N$, $k_{b2}$, and $k_{e}^{CD8}$ may have the greatest potential to improve complete response rates in CAR-T therapy by enhancing effector cell expansion and sustaining immune-mediated tumor control under parameter uncertainty.
@@
3.2 Model Validation
A set of 7500 virtual patients was generated using patient parameter values given in Kirouac et al. (2023); parameters were randomly chosen in the ranges given by patient data. Each patient was simulated for 1000 days under three possible scenarios: (1) no treatment, (2) treatment with CD4:CD8 ratio of 1:1, and (3) treatment with unspecified CD4:CD8 ratio (empty CD4+ compartment, as in the original Kirouac et al. model). In each case, patients were categorized into non-responder (NR) or complete responder (CR) depending on the tumor burden (greater or less than $10^{9}$ cells, respectively) at the end of the simulation time.

### v2

@@ Abstract
Chimeric antigen receptor (CAR) T cell therapy has shown remarkable success in hematological malignancies, yet patient responses remain highly variable and the roles of CD4+ and CD8+ subsets are not fully understood. We present an extended mathematical framework of CAR-T cell dynamics that explicitly models CD4+ helper and CD8+ cytotoxic lineages and their interactions with tumor antigen burden. Building on the Kirouac et al. (2023) model of antigen-regulated memory–effector–exhaustion transitions, our system of differential equations incorporates CD4+-mediated modulation of CD8+ proliferation, cytotoxicity, and memory regeneration through biologically grounded, saturating interactions. Sensitivity analyses identify effector proliferation, antigen turnover, and CD8+ expansion rates as dominant drivers of treatment outcome. Virtual patient simulations recover reported qualitative trends in CAR-T composition, including enhanced expansion and tumor clearance for defined CD4:CD8 products relative to CD8-only formulations, while also revealing inter-patient variability and time-dependent effects. To assess the practical limits of patient-level prediction under parameter uncertainty, we introduce controlled noise into key parameters and show that direct mechanistic classification rapidly degrades. We then demonstrate that a simple feed-forward neural network can partially recover predictive signal from noisy inputs, outperforming a naïve baseline while remaining consistent with mechanistic sensitivities. This work positions the extended model as a hypothesis generator, and illustrates how data-driven methods can complement mechanistic modeling when parameter uncertainty constrains predictive confidence.
@@ 1 Introduction
CD4+ and CD8+ T cells are distinct lineages of T cells (Germain et al. 2002), each with unique and complementary roles in the immune response. CD8+ T cells, known as "cytotoxic" T cells, specialize in directly killing foreign cells (Andersen et al. 2006). They are equipped with the ability to recognize antigens presented by infected or malignant cells and trigger apoptosis (cell death) in those cells. CD4+ T cells, often referred to as "helper" T cells, play a crucial role in orchestrating the immune response by secreting cytokines that activate other immune cells, including CD8+ T cells (Ahrends et al., 2013; Ahrends et al., 2017). Importantly, CD4+ and CD8+ T cells are separate lineages; they do not switch from one to the other, but instead, each lineage maintains its specialized function throughout its lifespan. In cancer immunotherapy, the roles of both CD4+ and CD8+ T cells are critical, as CD4+ cells provide necessary signals to optimize CD8+ T cell activity, creating a synergistic effect in tumor elimination (Boulch et al. 2021).
It has been shown that a 1:1 ratio of CD4+ to CD8+ CAR-T cells yields optimal therapeutic outcomes (as opposed to an unspecified ratio dependent on patient T cell composition), balancing efficacy and toxicity across multiple dosing levels (Turtle et al. 2016). Motivated by these findings, we extend the Kirouac et al. model to explicitly include CD4+ and CD8+ T cell lineages and their distinct functional roles. This formulation enables the simulation of various CD4:CD8 ratios and their impact on tumor–immune dynamics. We extend this framework by modeling separate CD4+ and CD8+ CAR-T cell lineages, thereby enabling the investigation of their distinct functional roles and synergistic interactions (Figure 1). Importantly, the model does not explicitly impose any optimal CD4:CD8 ratio; rather, we explore how effective tumor clearance may emerge from the complementary requirements of sufficient CD8+ cytotoxic effector activity and CD4+ helper-mediated support. In particular, the 1:1 ratio is examined because of its clinical relevance and established use in defined-composition CAR-T products.
@@
The goal of this work is therefore twofold. First, we extend the Kirouac et al. framework by explicitly separating CD4+ and CD8+ CAR-T cell populations in order to mechanistically examine how the ratio of these two lineages (i.e. the balance of helper and cytotoxic roles) may shape treatment outcomes. Second, we assess the extent to which such a mechanistic model can support patient-level outcome prediction in the presence of biological heterogeneity and measurement noise. Rather than performing a comprehensive bifurcation or dynamical systems analysis, our focus is on identifying which parameters most strongly influence clinically relevant outcomes, and evaluating how robust prediction is dependent on measurement of these parameters.
@@ 2 Methods
In this study, we develop a coupled mathematical model as a minimal extension of the CAR-T cell framework proposed by Kirouac et al. (2023). Tumor growth, antigen dynamics, and the antigen-regulated memory-effector-exhaustion structure of CAR-T cells are retained unchanged. The primary extension is the explicit separation of the CAR-T population into distinct CD4+ and CD8+ lineages, together with CD4+-mediated modulation of CD8+ cytotoxicity and memory regeneration. When the CD4+ dose is set to zero, all CD4+ compartments remain identically zero and the CD4-mediated modulation terms reduce to unity, such that the system collapses to the original Kirouac et al. formulation (up to relabeling of the aggregate CAR-T population). This structure enables systematic exploration of CD4:CD8 composition effects while preserving the original model as a limiting case. By modeling the interactions between these cell lineages and tumor cells, we aim to predict the therapeutic outcomes of CAR-T cell therapies with varying CD4:CD8 ratios. The model also allows for an exploration of the effects of patient-specific variations on treatment response. A schematic representation of the model is shown in Figure 1.
@@ 2.1 Tumor and Antigen Dynamics
$\small\frac{dB}{dt}$ $=\mu_{B}\left(1-\frac{B}{B_{\max}}\right)B-k_{\text{kill}}^{CD8}\left(1+\eta_{k}\,\frac{T_{E2}^{CD4}}{K_{\eta}+T_{E2}^{CD4}}\right)\,\left(\frac{{T_{E2}^{CD8}}^{k_{t}}}{TK50^{k_{t}}+{T_{E2}^{CD8}}^{k_{t}}}\right)B$
@@
where $\eta_{k}$ denotes the maximum fold-increase in CD8+ cytotoxicity attributable to CD4+ T cell-derived cytokines, and $K_{\eta}$ is the half-saturation constant (Ahrends et al. 2017). Biologically, this formulation reflects the role of CD4+ T cells, particularly Th1-like CAR-T cells, in secreting cytokines such as interferon-gamma (IFN-$\gamma$) and interleukin-2 (IL-2), which are known to enhance the activation, survival, and cytotoxic function of CD8+ T cells (Boulch et al. 2021). IFN-$\gamma$ also promotes tumor antigen presentation and induces tumor cell apoptosis; accordingly, the final term in Eq. (1) represents an effective CD4+-mediated contribution to tumor control through cytokine signaling rather than direct cytotoxic killing (Bove et al., 2023).
@@ 2.2 CD8+ CAR-T Cell Dynamics
The dynamics of CD8+ CAR-T cells are modeled by a set of differential equations that model the transitions between different T cell states: memory ($T_{M}^{CD8}$), early effector ($T_{E1}^{CD8}$), late effector ($T_{E2}^{CD8}$), and exhausted ($T_{X}^{CD8}$) (Kirouac et al. 2023). These equations take into account the proliferation, differentiation, death rates, and interactions with tumor antigens ($B_{A}$).
@@
$\small\frac{dT_{M}^{CD8}}{dt}$ $=\mu_{M}^{CD8}\left(2f_{\text{max}}^{CD8}\left(1-\frac{B_{A}^{k_{m}^{CD8}}}{B50^{CD8^{k_{m}}}+B_{A}^{k_{m}^{CD8}}}\right)-1\right)T_{M}^{CD8}$
@@ 2.3 CD4+ CAR-T Cell Dynamics
CD4+ T cells, or helper T cells, play an essential role in coordinating the immune response. Unlike CD8+ T cells, the main role of CD4+ cells is not directly killing infected or cancerous cells (although they can induce cell death by themselves) (Boulch et al. 2021). Instead, their main function involves enhancing the function of other immune cells, including CD8+ T cells, through the secretion of cytokines such as interferon-gamma (IFN-$\gamma$) and interleukin-2 (IL-2) (Ahrends et al. 2017). These cytokines stimulate the expansion and survival of CD8+ T cells, thereby augmenting their cytotoxic effects on tumor cells. In CAR-T therapy, CD4+ T cells are critical for sustaining an effective immune response, promoting memory formation, and improving tumor clearance (Boulch et al. 2021). CD4+ modulation of CD8+ function is represented using a saturating Hill-type term, reflecting cytokine-mediated helper effects that are nonlinear and saturable in nature. This choice provides a minimal and biologically grounded interaction without introducing additional state variables or exacerbating parameter identifiability issues.
@@
$\small\frac{dT_{M}^{CD4}}{dt}$ $=\mu_{M}^{CD4}\left(2f_{\text{max}}^{CD4}\left(1-\frac{B_{A}^{k_{m}^{CD4}}}{B50^{CD4^{k_{m}}}+B_{A}^{k_{m}^{CD4}}}\right)-1\right)T_{M}^{CD4}$
@@ 2.4 Model Parameters
A list of parameters used in the model is given in Table 1. Kirouac et al. reported physiologically feasible ranges and averages for the parameters in their original model, which did not have separate CD4+ and CD8+ compartments (Kirouac et al. 2023). However, several parameters are likely to differ biologically between these two lineages (Table 2).
@@
CD8+ CAR-T cells are primarily responsible for direct cytotoxicity and tend to differentiate more rapidly into effector cells ($\mu_{E}$) but are also more prone to exhaustion ($k_{ex}$) and activation-induced cell death ($d_{E}$, $d_{X}$) (Turtle et al. 2016; Hinrichs et al. 2011; Foulds et al. 2002; Fraietta et al. 2018; Gauthier et al. 2021; Blank et al. 2019; Siddiqui et al. 2019). In contrast, CD4+ CAR-T cells play a predominantly helper role through cytokine production (e.g., IFN-$\gamma$, IL-2), are generally more resistant to exhaustion, and may exhibit greater longevity and plasticity, potentially resulting in higher rates of reversion to memory ($r_{M}$). As such, when fitting the model to data or simulating lineage-specific behavior, these parameters should be relaxed and estimated separately for CD4+ and CD8+ populations. This refinement is supported by experimental evidence from preclinical and clinical studies, including those by Turtle et al. (2016), Fraietta et al. (2018), and Gauthier et al. (2021), which consistently demonstrate phenotypic and functional distinctions between CD4+ and CD8+ CAR-T cells in vivo.
@@ 2.5 Sensitivity Analysis
Because the Kirouac et al. framework is parameterized using high-dimensional, correlated distributions derived from patient data, isolating low-dimensional bifurcation structure is not straightforward. While nonlinear transitions in treatment outcome can be observed when varying individual parameters or parameter pairs, these thresholds are not fixed across patients and are strongly modulated by variability in other biological processes. For this reason, we focus on sensitivity-based analyses that better reflect clinically relevant heterogeneity.
There are two main types of sensitivity analysis: local and global. Local sensitivity analysis involves perturbing each parameter individually from a baseline value while holding other parameters constant, and looking at the change in output (Qian et al. 2020). In this context, we applied finite difference perturbations (1%) to each parameter (starting from baseline patient parameter values) and initial condition individually while tracking changes in the area under the curve (AUC) of tumor burden over 1 year. Sensitivities were computed as normalized relative changes, defined as the relative change in the output (area under the curve) normalized by the relative change in the input, allowing us to rank the most influential parameters under the baseline scenario.
@@ 3.1 Sensitivity Analysis
Local sensitivity analysis (Figure 2) revealed that the most sensitive parameters were the proliferation burst factor during antigen stimulation ($N$), tumor antigen generation and clearance rates ($k_{b1}$ and $k_{b2}$), CD8+ T cell expansion dynamics (including the effector proliferation rate $k_{e}^{CD8}$ and proliferation rate $\mu_{M}^{CD8}$), and tumor-related parameters such as the antigen half-max ($B_{50}^{CD8}$) (Figure 2). These parameters exhibited the highest relative sensitivities; that is, small changes in their values can lead to substantial differences in tumor burden dynamics within the model.
@@
Global sensitivity analysis (Figures 3, 4) found that the parameter representing the number of effector cell doublings ($N$) exhibits the highest sensitivity in our model, with both PRCC and Sobol analyses indicating it as the strongest driver of tumor burden dynamics. Parameters governing the dynamics of the antigen-bound complex ($k_{b1}$, $k_{b2}$) and the CD8+ effector proliferation rate ($k_{e}^{CD8}$) also showed high sensitivities. These findings align with prior analysis by Kirouac et al. of their original mechanistic framework (2023), which identified effector proliferation and memory regeneration as critical determinants of treatment response. Our analyses suggest that modulating parameters such as $N$, $k_{b2}$, and $k_{e}^{CD8}$ may have the greatest potential to improve complete response rates in CAR-T therapy by enhancing effector cell expansion and sustaining immune-mediated tumor control under parameter uncertainty.
@@
3.2 Qualitative consistency of model with reported CAR-T composition effects
Because suitable clinical datasets containing matched CD4:CD8 ratios, longitudinal CAR-T expansion, and outcome measures remain limited, we do not attempt a full quantitative validation of predicted response or survival. The original Kirouac et al. model has been previously analyzed against available clinical data; our extension builds on this framework by explicitly resolving CD4+ and CD8+ CAR-T cell lineages. Accordingly, this section focuses on whether reported qualitative trends in CD4:CD8 composition and treatment outcome can emerge from the extended model under biologically grounded assumptions, positioning the framework as a mechanistic hypothesis generator rather than a finalized clinical prediction tool.
A set of 7500 virtual patients was generated using patient parameter values given in Kirouac et al. (2023); parameters were randomly chosen in the ranges given by patient data. Each patient was simulated for 1000 days under three possible scenarios: (1) no treatment, (2) treatment with CD4:CD8 ratio of 1:1, and (3) treatment with unspecified CD4:CD8 ratio (empty CD4+ compartment, as in the original Kirouac et al. model). The no-treatment case serves as a control, the unspecified CD4:CD8 case corresponds to the original Kirouac et al. model with an empty CD4+ compartment, and the 1:1 CD4:CD8 treatment is included because of its clinical relevance, as previously described. In each case, patients were categorized into non-responder (NR) or complete responder (CR) depending on the tumor burden (greater or less than $10^{9}$ cells, respectively) at the end of the simulation time.
3.2.1 Consistency of model outcomes with reported effects of 1:1 CAR-T composition